UT 08-Jan-2020 - 2nd Half Eng

Plan

Night Center 00:26
  • DONE TCS 2020A - testing
    • Activation:
      • ssh tcs@tcs1
      • cd lbt/tcs
      • current/tcs-cluster switch 2020A
    • Reversion
      • current/tcs-cluster switch 2019B
    • Alternate method in case you fail to switch TCS version using "tcs-cluster", please switch manually:
      • exit all GUIs
      • as tcs@tcs2:
        • netconfig stop all
        • netconfig stop
      • as tcs@tcs1:
        • netconfig stop
        • cd /lbt/tcs
        • rm current
        • ln -s 2020A current
        • netconfig start
        • netconfig start all
      • start GUIs
    • Changes:
      • IIF
      • PCSGUI Mount/Rotator
        • shows rotator (azimuth and currently authorized left and right rotators for MODS and LUCI) orientation
        • visual clues are provided for positions < 0 and > 360 degrees
        • red line shows which limit (upper or lower) will be hit when tracking
  • DONE AGW-3 Checkout
    • DONE AGW-3 cutout/WFS issue - check on LUCI_R.cfg parameters
    • DONE LUCI-2 AGW-3 Rotator Zp - requires astrometric mask - seeing < 1.5"
    • LUCI-1/LUCI-2 bino Transform collection (requires 1.2-1.5" or better seeing on DIMM)
      • use -0.4 for LUCI-1
  • Poor Weather Backup
    • Local Mergers - any conditions up to 2.5" seeing
      • /home/observer/share/Service_Observing/Eng_Dec_2019/Poor_Weather_Backup/Local_Mergers - RM507-5/4
      • /home/observer/Service_Observing/Eng_Dec_2019/Poor_Weather_Backup/Local_Mergers - RM507-1-3
      • /home/lbto/Service_Observing/Eng_Dec_2019/Poor_Weather_Backup/Local_Mergers - OBS-N
      • Visible Targets (Priority Order):
        • UGC 4635 - 30min Bino, 1hr mono
        • NGC 3656- 30min Bino, 1hr mono
        • NGC 2655- 30min Bino, 1hr mono (best between 2-5am due to telluric delta airmass)
        • Arp 156 -30min Bino, 1hr mono (best between 2-5am due to telluric delta airmass)
        • NGC 3597 (low elevation - never goes above airmass 1.7 - good if running out of targets) - 30min Bino, 1hr mono
    • H-band Library - Thick clouds and/or seeing >2.5"
      • /home/observer/share/Service_Observing/Eng_Dec_2019/Poor_Weather_Backup/H_band_Library - RM507-5/4
      • /home/observer/Service_Observing/Eng_Dec_2019/Poor_Weather_Backup/H_band_Library - RM507-1-3
      • /home/lbto/Service_Observing/Eng_Dec_2019/Poor_Weather_Backup/H_band_Library - OBS-N
        • HD 42543 M2I H=1.2
        • HD 44536 M2I H=0.7
        • HD 20797 M0Iab H=0.6
        • BD-15 1319 K5III H=6.4
      • ~30 min each including Telluric
    • ICRF3 USNO followup - flexible conditions - 25min per target Bino-LUCI (zJ LUCI-1 HK LUCI-2) -
      • Goal is to obtain spectroscopic redshifts for a new ICRF catalog (verify they are AGNs and not MW stars)
      • Can be done in any conditions
      • Designed for fraternal zJ(L1) and HK(L2) G200+1"slit, but LUCI-1 zJ only is sufficient backup
      • If one can see at least 2 bright emission lines - then object is complete. If conditions are really poor, take extra exposures
      • Scriptss are located in:
        • /home/observer/share/Service_Observing/Eng_Jan_2020/Backup/ICRF3_Rothberg_LD_2019B_006 - Rm507-5/4
        • /home/observer/Service_Observing/Eng_Jan_2020/backup/ICRF3_Rothberg_LD_2019B_006 - Rm507 1-3
        • /home/lbto/Service_Observing/Eng_Jan_2020/backup/ICRF3_Rothberg_LD_2019B_006 - OBS-N
          • finder charts are in J-band from 2MASS images
          • OT file: ICRF3_LUCI.xml
          • 26 targets spread across all of the sky
          • Following are visible (priority order)
            • 0426+273
            • 0614+570
            • 0854-108
            • 0509+293
            • 0509+406
            • 0513+276
            • 0535+001
            • 0600-000
            • 0559+422
            • 0600+063
            • 0637-128
            • 0641-135
            • 0702+612
            • 0804-055
            • 0845-068

Summary

AGw3: The preparatory work that John & Barry did and the changes John made earlier today seemed to have helped a lot with the cutout problem. We were able to move on to, for both sides: 1/ checking the rotator zeropoints; 2/ collecting data to measure the rotator center; and 3/ collecting data for the AGw-PCS transform. Seeing was for the most part steady <~0.8", however there were some excursions to 1.5". John noted that on DX there was some trefoil - Z9 was ~ -400nm, but otherwise collimation looked good.

LUCI
  • There was a LUCI2 MOS error for which I had to call Dave. The problem was that the translator did not quite reach the position of the N30 FS mask in storage. It was one or two steps out of tolerance.
  • It looks like there is a ~210 pix shift in the wavelength calibration between the two LUCIs for G210 @ 2.3 mic. The same sky line appears at a position ~210 pix to the left on L2 than where it falls on L1.
We switched to TCS build 2020A before reconfiguring at the turnover between LBTI & LBTO eng. We had some issues running the commands provided in the email, but did get 2020A working. Steve has written more details on this.

We ended the night with a science target. For tomorrow, it is unlikely that the transform data will be analyzed and unless there are pressing engineering tasks which I'm not thinking about right now, we perhaps could run D or other science.

Preparations

04:30 UT JMH updated LUCI_R.cfg with the following THREE changes (see Com20200107 for analysis of last night's data).
  • WFS_pupil_shift_y changed from -1.0 to +1.0
  • WFS_pupil_dir_x changed from +1 to -1
  • WFS_pupil_dir_y changed from -1 to +1

Details

07:20 - noted LUCI2 Field a bit offcenter - 20-30pix strip at left edge. recentered by adjusting about 20 pix. LUCI1 was OK.

07:23 - handing over

LBTI de-Souled DX and left SX as is (SL)

TCS Build 2020A

07:25 Steve switched to TCS build 2020A

- problem running tcs-cluster switch 2020A command -- complains that cannot kill GUI, pidof command not found - Steve following the instructions for "manual" -- could not stop ECS... -- not going too smoothly

07:43 Reconn to DD, started GCSGUI left/right and IIFGUI

07:53 - New build, new TCS config files needed (LUCI2 is trying to select AGW2).

John is editing files in /lbt/tcs/current/gcs/configuration/GCS (as noted in IT7252) after which Steve restarted GCSR.

07:59 Going to a bright star on meridian and in south.
  • RPC error: Wed Jan 8 08:02:16.485 2020 TCS authorize alert for right LUCIFER at bentGregorianFront focus: Invalid buffer: RPCERROR connectStatus. Steve restarted AOS which fixed the problem.
  • Paddle moves backwards in x on AGw3 because of flip in readout.

Checking GCS WFS parameters for AGW3@RFBG

08:14 0,565.0 Pointing & Collimation check on slightly off-axis star. PA=158. wfs spots slightly low. Looks good, collimation reached.

08:15 0,612.4 On-axis preset. AGW3 Pupil is sort of near the right edge on GCSR, but it is perhaps equally near the left edge on GCSL.

08:20 24,571.6 Same off-axis guide star as before (08:14) but with a different PA, 30 deg off of what it was earlier, PA=128.
  • Pupil is now towards the bottom of the cutout... alas, more changes needed. wfs spots are a full row too low.

08:27 41,589 New preset: same off-axis guide star but another 30 deg PA change, now PA=98.
  • Pupil is now sort of like it was on-axis. Up-down pretty well-centered.

08:30 0,455 New preset: this time guide star far off-axis. R=262". Pupil looks good.

Checking Rotator Zeropoints (both sides)

08:33 Rotator Zero Points with Stone E astrometric mask
  • airmass = 1.3, some vignetting from tertiary. Cutout issue looks good.
  • L1 images FWHM ~4pix, L2 images also ~4pix
  • L1: 2,3,4 and L2: 4,5,6
  • Mask-to-detector:
    • L1: X-shift, Y-shift = -3.94, 14.71 pix and angle 0.1827.
    • L2: X-shift, Y-shift = -9.00, -3.9 pix and angle -0.2673
  • Rot Matrix:
    • L2: dX,dY,dtheta = -0.52", -0.93" -0.09 deg after removing stars 31,32,33.
    • L1: dX,dY,dtheta = 3.2", -1.7", -0.02 deg after removing stars 31,32.
    • L2: dX,dY,dtheta = -0.52", -0.81, -0.08 deg after fixing star 32 and deleting stars 30,31,33. * Sending offsets and taking confirm images:
    • L1: 5 - stars look pretty well centered in boxes (5 minus 2 - slow sub, high cpu load on obs3 noted yesterday and today, 29 min ago today, well before sub started but while using LUCI). RTD wonky. Running it from obs panel on obs3 and it 1/ took very long time to do subtraction and 2/ subtracted pair look offset.
    • L2: 7 - same - stars look pretty well centered in boxes.

Conclusion: Rotator zeropoints that are currently in place look good for both L1 and L2. These are, for L1/LFBG, -187.735, and for L2/RFBG, +176.85.

09:15 Stone E is already setting (AM~1.34 now), so Stone I for transform data collection... But, check Rotator Center on Stone E.
  • John notes that we're accumulating Z9 on DX active optics.

09:15 Active preset PA=0 to dense subregion of Stone E (near star 54). -12.2,472.7

09:20 Opening LUCI on rm507-5 (Fedora) now. OK, but it shows me luci2.20200108.0008 as an N3.75 no mask image where as on the obs3 rtd it correctly shows the N30 FS.

09:28 Rotator Center images, all track mode (zen-corr DIMM ~ 0.6")
  • PA=0 track L1 7 and L2 9
  • PA=180 track L1 8 L2 10
  • PA=90 track L1 9 L2 11
  • PA=270 track L1 10 L2 12
  • darks just in case L1 11 L2 13

09:45 Preset to Stone I field for a second mask alignment. +58.5,491.5 (spots 1 row low)
  • L1 images: 12,13,14 and L2 images: 14,15,16
    • L2 images look a pretty elongated, FWHM ~ 8,5.3" but L1 images ~round, FWHM ~ 6.8, 5.7
  • Mask-to-det (X-shift,Y-shift,Angle):
    • L1 (-8.242, 17.9591, 0.1783)
    • L2 (-9.5424, -1.8131, -0.2633)
      • Note: LMS file was found for L1 but not for L2, had to copy it over to rm507-5 machine.
  • Rot Matrix
    • L1 (2.27", -0.3966", 0.0307 deg) using all stars
    • L2 (-1.26", -1.37", -0.109 deg) using all stars but 35. Erring in same direction - marginal, but John says to take transform data.
  • Confirmatory images:
    • L1 15 and L2 17 look good.

09:52 Refining WFScam_hotspot_y 237.6 ---> 230.6

09:58 Refining WFScam_hotspot_sext_x 60.0 ---> 61.5

-400 Z9 seems to be irritated by a bad S-H spot just lower right of center

StoneI Transform Data Collection (both sides)

10:15 Send active preset to Stone I GS=TS=2 (14<mag<16) and PA=-64
  • Use modified list at PA=-64 and PA=-90. It is not a very dense field.
  • Putting N30 FS mask in place

10:17 on-axis preset (spots well centered, Z9 decreasing from -400 to -300 nm)

10:26 Refining guidecam_hotspot_y 215.0 ---> 217.0, guidecam_hotspot_x 277.5 ---> 276.5

10:28 L2 mask error in moving from storage to turnout. Called Dave. Dave resolved the problem. The translator failed to reach the position of the N30 Field Stop after putting away the previous mask; it was 1 step out of tolerance.

10:55 Test H-band images on on-axis star. 5-sec exposure is good, giving about 20,000 counts in the peak. There are 2 other faint stars in the field - one near the top edge and the other not near an edge but fainter.

11:02 transform_collect started.
  • problem with left side: Error in station_check: more than one focal stations matches request      station_params.agw_name = AOAGW1
    • John knows where to fix this. Removed AOAGW1 from the LARGOS entry in station_define.pro. Restarting after the fix.

L1 L2 GS comment
19 21 on-axis image  
20 22 background  
21 23 14 seeing blew up to 1.5 during 23, delay in taking L1 - 21 taken during good seeing
22 24 9 seeing 1.4 during 24
23 25 10 seeing back to 0.8
24 26 3 GCSL oscillated back and forth 11:19, also close enough to be affected by guide probe shadow
25 27 13  
26 28 1  
27 29 5 27 at 11:27 UT
28 30 0  
29 31 12  
30 32 11 seeing blew up to 1.5"
31 33 15  
32 34 4 strong gradient on 34, could be corner of guide probe shadow
33 35 6 L1, GCSL took a while, looked like it was going to start osc again
34 36 16  
35 37 10 now start PA=-90
36 38 3  
37 39 13  
38 40 12  
39 41 11 GCSR osc 1 cycled before settling
40 42 15 GCSL failed to get GS=15, Steve tweaked pointing, restarted script and took image 40.
41 43 4 seeing 0.4" on GCSR, gradient top-to-bottom due to probe shadow
42 44 6  
43 45 8  
44 46 16  
45 47 20 neither left nor right could not reach 20*
45 47 6  
46 48 8  
47 49 16  
48 50 on-axis  

* GCSR said it looked to be right at the edge (x,y): 175.727, 11.7759 mm) | and GCSL: 175.731, 11.7551 mm.

11:19 Steve restarted GCSL after oscillation problem.

On left side, transform_collect, exp=5.0, start=4. Looks solid now. L1: 24 is star 3

Around 11:50. On right side, star 20 is out of range. It is probably just out of the patrol field - to far to the bottom right.

11:51: On left side, star 15 at PA=-90 (star 20) was not found. The star was not in the acq image. Steve tweaking pointing. After that, transform_collect, exp=5.0, start=20.

11:59: Restart right side at star 22 (GS 6) to let it finish nicely with the on-axis image. (It had crashed out).

~12UT: Transform data collection complete. In PCSInstrument.conf, LUCIFER LEFT RZP = -187.735 and LUCIFER RIGHT RZP = +176.85. Will need to take a couple of DIT= 5-sec darks (Done at the end of the night - these files are L1: 63-65 and L2: 67-69).

12:20: Slewing to telluric for UGC4635. GCSR looks a bit pincushion-like to start.
  • Telluric spectra: L1 52-54 and L2 54-56
    • NOTE: using same config on L1 and L2 (G210 @ 2.3 mic), there is a wavelength shift between the two sides (spectral range for L2 is redder).
  • AM = 1.73
12:44 Slewing to object, UGC4635.
  • AM = 1.4 at start
  • The x2go session from rm507-3 to obs3 crashed just as I was centering up on slit. Moved to acquire on rm507-5.
    • Confirm images were L1 58 and L2 60.
      • L1 58 looks pretty good.
      • Used telsvc to offset right side by 0.24" and now taking L2 61. Too far to right. Offset -0.2" now. L2 62 ...tad left but pretty good.
  • Starting science as we are into twilight now. Sun is about -15 deg below horizon.
    • AM=1.52-1.73
    • Took 3 600-sec exposures. L1: 59-61 and L2: 63-65.
    • Let it go onto a 4th, but it is getting light and WFS may not hold on. With about 20-sec left, lost WFS.
13:51 Closing up

13:55 Taking 5-sec darks for transform data. These will be L1: 63-65 and L2: 67-69

14:05 Taking calibrations for UGC4635 science. These are L1: 66-83 and L2: 70-87.

14:30 Both LUCIs are in the end of night configuration. Goodnight...

-- BarryRothberg - 07 Jan 2020
Topic revision: r9 - 09 Jan 2020, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback